Question 5 (60 points) For the following first order ODE: 1) xy + (3x - 2y = 0, Xo = 0 DO 3X2 @O 2) (1-x)y + 2y = 0, xo = 0 please solve with power series. part b for 2and 3 is what im really struggling with
Answer #1

- for differial equation y + b(x) y = 0 0 A point, a, is point of singularity if p(x) is NOT continuous at ä. ats - o aso qboint solution - 16.) : x=0 is NOT the singular - ODE has a power series - at No =0. (i) azt is the only singular i R = Mir ļ- but this in given ope. c ² Cn och =0 = 0 xn- en i nGXn-1 + 2 (2-x) n0 n=1 .00, +2san x-0 >> Žnen an! - na an-1+1 n=0 n=1 &nxn- n=1 - no - 3 6 xeno - (n -2) (n 2 30 - In est sum of 3 replace n by nth E Here nit to ao l earlies) n+:1 to ao Cafter rnad 6 = ģu - * (-2)6 – Co n=2 Cz - og lg = 0 nas CH = C2 = 0 (as (p=0) in Cq=0=4=G=-.. Now as y= 2 e conoch isy - Có +612 + (